4. A drug manufacturer claims that fewer than 10% of patients who take its new drug for treating Alzheimer’s disease will experience nausea. In a random sample of 250 patients, 23 experienced nausea. Perform a significance test at the 5% significance level to test this claim.
5. A manufacturer of sprinkler systems designed for fire protection claims that the average activating temperature is at least 135. To test this claim, you randomly select a sample of 32 systems and find the °Fmean activation temperature to be 133◦F with a standard deviation of 3.3 ◦F. At α = 0.10, do you have °Fenough evidence to reject the manufacturer’s claim? Let a= .0.05
1
Expert's answer
2019-11-14T09:52:26-0500
  1. Given that
"p=10\%=0.1,x=23,n=250"

Determine the hypotheses

"H_0:p=10\%=0.1"

"H_1:p<0.1"

Is the sample large enough?


"np=250(0.1)=25\geq15""n(1-p)=250(1-0.1)=225\geq15"

Test statistic

The sample proportion is the number of successes divided by the sample size


"\hat{x}={x \over n}={23\over 250}=0.092"

Determine the value of the z-statistic:


"z={\hat{p}-p \over \sqrt{{p(1-p) \over n}}}""z={0.092-0.1 \over \sqrt{{0.1(1-0.1) \over 250}}}\approx-0.422"

The significance level is "\alpha=0.05," and the critical value for a left-tailed test is "z_c=-1.64"

The rejection region for this left-tailed test is "R=\{{z:z<-1.64}\}"

Since it is observed that "z=-0.422\geq-1.64=z_c," it is then concluded that the null hypothesis is not rejected.

Determine the "p-" value


"p=P(Z<-0.421637)\approx0.336645"

Since "p=0.336645\geq0.05," it is concluded that the null hypothesis is not rejected.

Therefore, there is not sufficient evidence to indicate that fewer than 10% of patients who take its new drug for treating for treating Alzheimer’s disease will experience nausea, at the 0.05 significance level.


5. The provided sample mean is "\bar{X}=133" and the known population standard deviation is "\sigma=3.3," and the sample size

is "n=32."

The following null and alternative hypotheses need to be tested:

"H_0: \mu\geq135"

"H_1:\mu<135"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.1," and the critical value for a left-tailed test is

"z_c=-1.28"

The rejection region for this left-tailed test is "R=\{{z:z<-1.28}\}"

The z-statistic is computed as follows:


"z={\bar{X}-\mu \over \sigma/\sqrt{n}}"

"z={133-135 \over 3.3/\sqrt{32}}\approx-3.4284"

Since it is observed that "z=-3.4284<-1.28=z_c," it is then concluded that the null hypothesis is rejected.

The p-value is


"p=P(Z<-3.4284)\approx0.0003"

Since "p=0.0003<0.1," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 135, at the 0.1 significance level.

There is enough evidence to reject the manufacturer’s claim, at the 0.1 significance level.


Let the significance level is "\alpha=0.05" The critical value for a left-tailed test is "z_c=-1.64"

The rejection region for this left-tailed test is "R=\{{z:z<-1.64}\}"

The z-statistic is "z=-3.4284"

Since it is observed that "z=-3.4284<-1.64=z_c," it is then concluded that the null hypothesis is rejected.

The p-value is


"p=P(Z<-3.4284)\approx0.0003"

Since "p=0.0003<0.05," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 135, at the 0.05 significance level.

There is enough evidence to reject the manufacturer’s claim, at the 0.05 significance level.